# Direct-to-Cell: A First Look into Starlink’s Direct Satellite-to-Device Radio Access Network through Crowdsourced Measurements

### v1

@@ IV-D Performance assessment of future DS2D services
As noted above, the FCC authorized a $10$-dB increase in OOBE for DS2D transmissions in March 2025. Although measurements from April 2025 do not yet show any improvement in SINR or other metrics, this regulatory change could enable SINR gains of up to $10$ dB in the near future, provided that transmitter non-linearities do not necessitate an increased power back-off with the new OOBE limit, and that the increased radiated power does not lead to significantly higher interference, which would partially offset the SINR improvement. These assumptions are reasonable, as such effects are expected to be minimal in the current DS2D deployment, where transmit power is limited by regulatory limits and performance is constrained more by received power and noise than by interference. Assuming a full $10$-dB SINR gain, the estimated spectral efficiency rises to $2.35$ (mean) and $2.38$ bps/Hz (median)—a threefold increase which results in around $11.7$ Mbps. As expected, this aligns with the well-established principle that increasing SINR yields significant capacity gains in coverage-limited scenarios.
@@
In Subsection III-C, a maximum MCS has been assumed corresponding to $256$-QAM. The 3GPP standard contemplates the use of even higher order modulation. However, this will not produce a noticeable capacity improvement. The reason is, again, the coverage-limited nature of the network: large values of SINR occur very infrequently with the existing SINR distribution, and thus the higher MCS are hardly ever used. This is the case even if the previously discussed $10$-dB increase is incorporated.
@@
Throughput in the DS2D network may also be increased by expanding the available bandwidth. Assuming that the additional spectrum is paired with proportional transmit power, the spectral efficiency remains constant, and the per-beam capacity doubles. Starlink could pursue this by accessing more mobile spectrum (IMT) for the SCS services, but due to the risk of interference, this would require terrestrial operators to release spectrum, which is unlikely. Alternatively, Starlink may seek to use Mobile Satellite Service (MSS) spectrum and has, in fact, requested access to several bands around 2 GHz (specifically, the 1.6/2.4 GHz bands and the 2020–2025 MHz Earth-to-space band). However, this request was denied by the FCC in March 2024.444See FCC DA-24-300 In an additional $2\times 5$ MHz were granted in the future, the estimated mean throughput could increase from $4$ Mbps to $8$ Mbps and could potentially reach values close to $24$ Mbps when combined with gains from the recent increase in the OOBE limit.
@@
In the longer term, satellite swarm array antennas show promising preliminary results in increasing the directivity of satellite beams [15], which allows for higher directivity gains and improved spatial multiplexing. Achieving this would require the launch and operation of many more satellites, which would also increase the overall throughput of DS2D networks and the effective capacity experienced by end users. Although the FCC denied Starlink’s request to authorize 30,000 DS2D-capable satellites, the company is still authorized to operate up to 7,500 such satellites, indicating substantial remaining potential for network densification.

### v2

@@ IV-D Performance assessment of future DS2D services
As noted above, the FCC authorized a $10$-dB increase in OOBE for DS2D transmissions in March 2025. Although measurements from April 2025 do not yet show any improvement in SINR or other metrics, this regulatory change could enable SINR gains in the near future. The expected gain is less than $10$ dB, because, on one hand, transmitter non-linearities require an increased power back-off with the new OOBE limit; and, on the other hand, the increased radiated can lead to significantly higher interference. The latter effect is expected to be small, as currently performance is mainly constrained by received power and noise rather than by interference. As for the former, SpaceX estimates [^4] See https://www.fcc.gov/ecfs/document/10914226871971/1 that the increased OOBE limit would improve SINR by $3.1$ dB. Assuming this gain, the estimated spectral efficiency rises to $1.17$ (mean) and $1.05$ bps/Hz (median), i.e. it improves by a $1.5$ factor, which results in around $5.9$ Mbps. As expected, this aligns with the well-established principle that increasing SINR yields significant capacity gains in coverage-limited scenarios.
@@
In Subsection III-C, a maximum MCS has been assumed corresponding to $256$-QAM. The 3GPP standard contemplates the use of even higher order modulation. However, this will not produce a noticeable capacity improvement. The reason is, again, the coverage-limited nature of the network: large values of SINR occur very infrequently with the existing SINR distribution, and thus high MCS are hardly ever used. This is the case even if the previously discussed increase in the OOBE limit is incorporated.
@@
Throughput in the DS2D network may also be increased by expanding the available bandwidth. Assuming that the additional spectrum is paired with proportional transmit power, the spectral efficiency remains constant, and the per-beam capacity doubles. Starlink could pursue this by accessing more mobile spectrum (IMT) for the SCS services, but due to the risk of interference, this would require terrestrial operators to release spectrum, which is unlikely. Alternatively, Starlink may seek to use Mobile Satellite Service (MSS) spectrum and has, in fact, requested access to several bands around 2 GHz (specifically, the 1.6/2.4 GHz bands and the 2020–2025 MHz Earth-to-space band). However, this request was denied by the FCC in March 2024.555See FCC DA-24-300 If an additional $2\times 5$ MHz were granted in the future, the estimated mean throughput could increase from $4$ Mbps to $8$ Mbps and could potentially reach values close to $12$ Mbps when combined with the increase in the OOBE limit.
@@
In the longer term, satellite swarm array antennas show promising preliminary results in increasing the directivity of satellite beams [15], which allows for higher gains and improved spatial multiplexing. Achieving this would require the launch and operation of many more satellites, which would also increase the overall throughput of DS2D networks and the effective capacity experienced by end users. Although the FCC denied Starlink’s request to authorize 30,000 DS2D-capable satellites, the company is still authorized to operate up to 7,500 such satellites, indicating substantial remaining potential for network densification.